translate to a algebratic equation then solve.
the perimeter of a rectangular carpet is 70 feet. the width is three-fourths the lenght. find the width.

Answer by RAY100(1637)

Perimeter = 2(Length + Width)
.
but per = 70,,and W=3/4 L,,,,,substituting
.
70 = 2(L + .75 L)=2(1.75L) = 3.5L
.
or L= 20,,,,,,W=.75*20=15,,,,feet
.
.
check,, p=2(L+W) =2(20+15)=2*35=70,,,,,ok
.